### v1

# Focus on Your Question! Interpreting and Mitigating Toxic CoT Problems in Commonsense Reasoning

Abstract
Large language models exhibit high-level commonsense reasoning abilities, especially with enhancement methods like Chain-of-Thought (CoT). However, we find these CoT-like methods lead to a considerable number of originally correct answers turning wrong, which we define as the Toxic CoT problem. To interpret and mitigate this problem, we first utilize attribution tracing and causal tracing methods to probe the internal working mechanism of the LLM during CoT reasoning. Through comparisons, we prove that the model exhibits information loss from the question in the shallow attention layers when generating rationales or answers. Based on the probing results, we design a novel method called $\mathbb{RIDERS}$ (Residual decodIng and sERial-position Swap), which compensates for the information deficit in the model from both decoding and serial-position perspectives. Through extensive experiments on multiple commonsense reasoning benchmarks, we validate that this method not only significantly eliminates Toxic CoT problems (decreased by 23.6%), but also effectively improves the model’s overall commonsense reasoning performance (increased by 5.5%).
Focus on Your Question! Interpreting and Mitigating Toxic CoT Problems in Commonsense Reasoning
Jiachun Li 1,2, Pengfei Cao 1,2, Chenhao Wang 1,2, Zhuoran Jin 1,2, Yubo Chen 1,2 Daojian Zeng 3, Kang Liu 1,2, Jun Zhao 1,2 1The Laboratory of Cognition and Decision Intelligence for Complex Systems, Institute of Automation, Chinese Academy of Sciences 2School of Artificial Intelligence, University of Chinese Academy of Sciences 3Hunan Normal University {jiachun.li, pengfei.cao, chenhao.wang, zhuoran.jin, yubo.chen, kliu, jzhao} @nlpr.ia.ac.cn zengdj916@163.com
1 Introduction
With the increase in scale, large language models (LLMs) have demonstrated outstanding performance in different reasoning tasks, among them, commonsense reasoning has received significant attention due to its importance for general intelligence. In this task, researchers have proposed a series of chain-of-thought (CoT) like techniques to elicit models’ potential abilities (e.g. Self-Consistency (Wang et al., 2023c), Least-to-Most (Zhou et al., 2023), Reflexion (Shinn et al., 2023)). Through them, LLMs can generate reasonable rationales and improve their reasoning performance.
Figure 1: Two examples for the Toxic CoT problem.
While these works have made great progress, we notice an overlooked problem in them, which we define as Toxic CoT — Sometimes LLMs can directly provide correct answers to questions, but after applying CoT-like methods, it brings extra reasoning paths to models, causing their answers to be wrong. Figure 1 illustrates two main error types of this problem — Rationale Drift and Answer Drift. Specifically, for the Rationale Drift case, given the question “What kind of status is the bald eagle given?”, the model can directly give the correct answer “protection”. However, in the rationale, the model explains “what is the bald eagle” as “a symbol of America”, which has a semantic drift from the question. Thus, the model chooses the wrong option “america” based on the drifting rationale. For the Answer Drift case, given the question “Metal is used to make what?”, the model can directly answer “instruments”. It can also generate a correct rationale “metal is to make tools and machines”, but when answering based on the rationale, the model drifts from it and selects the incorrect option “(4)”. We further conduct a statistical analysis over extensive commonsense reasoning datasets and find that, among all CoT errors, this problem accounts for 37% for the white-box model and 33% for the black-box model on average, indicating this problem has become a crucial bottleneck in CoT reasoning. [^1] Appendix A presents the details of settings and results in this statistical experiment.
So what is the mechanism behind this issue? In this paper, we attempt to answer this question by probing the inner workings of the LLM’s CoT reasoning. Specifically, we first make initial observations on examples of Rationale Drift and Answer Drift issues, which suggest that the model likely misses some important information from the question when generating corresponding rationales or answers. To further verify these findings, we use attribution tracing and causal tracing methods to probe the LLM in two stages (rationale generation stage and answer generation stage). By employing these methods under various experimental settings, we find that there is a significant loss of information flow from the question in the shallow attention layers when generating drifting rationales and answers. Therefore, we interpret the Toxic CoT problem as the model lacking information from the question in the two stages.
To validate our interpretation and mitigate this problem, we design an approach called $\mathbb{RIDERS}$ (Residual decodIng and sERial-position Swap) based on the interpretation. Concretely, for the Rationale Drift issue, we devise a decoding algorithm, promoting the model to generate tokens that pay more attention to question contexts. For the Answer Drift issue, we swap the positions of the output sequence, reducing the information loss from the question to the final prediction. We evaluate our method on five commonsense reasoning benchmarks and conduct extensive experiments. The results not only prove our interpretation, but also indicate that our method is effective in addressing the Toxic CoT problem and improving the model’s overall commonsense reasoning abilities.
We summarize the contribution of this paper as follows:
(1) We identify a crucial bottleneck affecting LLM’s reasoning performance called the Toxic CoT problem, probe this issue through attribution tracing and causal tracing methods, and interpret the mechanism behind it as the model missing information from questions in shallow attention layers. The results contribute to a more in-depth understanding of the LLM’s reasoning mechanisms.
(2) To mitigate the Toxic CoT problem, we introduce $\mathbb{RIDERS}$, which effectively compensates for the internal information loss during CoT reasoning from decoding and serial-position perspectives.
(3) We conduct extensive experiments on various benchmarks. The results not only verify the rationality of our interpretation, but also demonstrate the effectiveness of our method in addressing the Toxic CoT problem (the proportion of the problem decreased by 23.6%) and enhancing commonsense reasoning performance (overall accuracy increased by 5.5%). We will release the source code for this work if it is accepted.
2 Problem Statement
2.1 Toxic Chain of Thought Reasoning
We start our work by formally defining the Toxic CoT problem as follows:  [^2] In practice, CoT-type methods all have Toxic CoT problems, but to simplify the work, this paper mainly focuses on the basic CoT prompting.
Definition 2.1 (Toxic CoT).
Given a question $q$ and the correct answer $o^{*}$, if the model’s output $\mathcal{M}$ meets the following conditions, it is considered a case of Toxic CoT:
$o^{*}$ $=\mathcal{M}(q,P_{d})\wedge o^{*}\neq\mathcal{M}(q,P_{c})$
where $o^{*}=\mathcal{M}(q,P_{d})$ indicates the model’s direct answering for $q$ is correct, $o^{*}\neq\mathcal{M}(q,P_{c})$ indicates the model’s cot-like answering for $q$ is wrong, $P_{d},P_{c}$ are the corresponding prompts.
2.2 Two-stage Drift Issues
To investigate the reasons for the problem, we classify these Toxic CoT cases and identify a main error causing this problem (On average, it accounts for 67% on two datasets, see more details in Appendix B). Furthermore, if we divide the CoT process into two stages: rationale generation and answer generation, there exist two types of issues in this error:
Definition 2.2 (Rationale Drift).
If the reasoning chain is factually correct but logically inconsistent with the question, this case is called “Rationale Drift” (see Figure 1a).
Definition 2.3 (Answer Drift).
If the reasoning chain is both factually correct and logically consistent with the question, but the final answer is inconsistent with the rationale, this case is called “Answer Drift” (see Figure 1b).
2.3 Hypothesis Formulation
To provide a direction for subsequent probing experiments, here we attempt to propose hypotheses for the mechanism of issues by analyzing some examples. For the Rationale Drift issue, the model tends to focus on part of the essential reasoning conditions in the question context. As an example, in Figure 1a, the CoT only focuses on the “bald eagle” in the question but misses another key information “status”. As for the Answer Drift issue, the model seems to be disrupted by CoT, losing attention to the question and resulting in an off-topic prediction. For instance, in Figure 1b, though the CoT gives correct information “to make tools and machines”, the model can only predict the wrong answer “(4) metal fabrication shop”. This is likely because the model loses the question’s target “to make what” and directly copies the entity “metal”, which frequently appears in CoT, as the answer. Therefore, we summarize our hypotheses as follows:
Hypothesis 1.
The Rationale Drift issue arises from the model lacking information from the question context in the rationale generation stage.
Hypothesis 2.
The Answer Drift issue arises from the model lacking information from the question in the answer generation stage.
To validate the above hypotheses, in the following two sections, we conduct probing experiments, exploring the LLM’s internal working mechanisms during the two stages of CoT reasoning.
3 Tracing Information Flow in Rationale
[…]
3.1 Tracing Method
To investigate the roles of different model components in the rationale generation stage, we use attribution scores (Hao et al., 2021; Dai et al., 2022; Wang et al., 2023b) to compute the contribution of a neuron $\omega$:
[…]
where $F(\cdot)$ represents the model’s output. We compute the attribution score via Riemman approximation of the integration and $m$ is the number of approximation steps. For neurons $A^{(l)}$ in the i-th attention layer, we sum the absolute values of scores on all attention heads to get the final attribution score. Since the attention module involves interactions between different tokens, we can compute the information flow between the question context $q$ and the CoT $c$ on it:
$Q^{(l)}_{qc}=\frac{1}{|N|}\sum_{(i,j)\in C_{qc}}Attr(A^{(l)}_{i,j})$ (2)
$C_{qc}=\{(i,j)|q_{s}\leq i\leq q_{e},\ c_{s}\leq j\leq c_{e}\}$
Here, $Attr(A^{(l)}_{i,j})$ represents the intensity of information flow from the $i$-th token to the $j$-th token in the $l$-th attention layer and $|N|$ denotes the number of CoT steps. More implementation details of this method are reported in Appendix C.
(a) Llama2-13B
(b) Baichuan2-13B
Figure 2: Attribution tracing results on Winogrande.
3.2 Attribution Tracing Experiment
Experimental Settings
To validate the deficiency in Hypothesis 1, we need to figure out the context information flow difference between generating a drifting rationale and a correct one. Thus, we first use golden labels as hints to generate correct CoTs in the drifting cases. Then, we compute the average information flow under these two cases and compare their results. We choose Llama2-13B (Touvron et al., 2023) and Baichuan2-13B (Yang et al., 2023) as our probing models, since they are moderated-sized white-box models with decent CoT performance. For datasets, we select Winogrande (Sakaguchi et al., 2020) and CSQA (Talmor et al., 2019). [^3] Unless otherwise specified, we use the same models and datasets in the following probing experiments. The detailed implementation of this experiment is shown in Appendix C.
Result and Analysis
Figure 2 illustrates our experimental results on Winogrande (The results on CSQA are shown in Appendix C). We can find that: (Claim 1) When the Rationale Drift issue occurs, CoT receives less information from the question context compared to the correct case. On both datasets and different models, there is a significantly lower information flow between the question context and CoT when the LLM generates a drifting rationale (the blue line) compared to the correct one (the orange line). This aligns with Hypothesis 1. (Claim 2) The shallow attention layers are crucial for LLMs to extract contextual information. In all cases, both the information flow and the gap peak at around the 15th attention layer, indicating these layers are significant sites for the rationale generation.
(a) Llama2-13B
(b) Baichuan2-13B
Figure 3: Information flow divergence comparison on Winogrande.
Supplementary Experiment
In the main experiment, we use golden labels to generate the correct CoT. To eliminate the influence of this additional factor on our results, we design a supplementary experiment. Concretely, we first use the label to generate CoTs from correct reasoning cases, which serves as a control group. Then, we compute the context information divergence between the newly generated CoT $c_{n}$ and the original one $c_{o}$, i.e.:
$Attr(c_{n}|c_{o})=Q^{(l)}_{qc_{n}}-Q^{(l)}_{qc_{o}}$ (3)
where $q$ is the question context. We compare this divergence between the control and drifting group, whose results are reported in Figure 3 and Figure 10. As we can see, the gap between the correct CoT and the drifting one (the blue line) is larger than the control group (the orange line) and the max divergence occurs in shallow layers (around the 15th layer). This indicates that a correct CoT indeed gets more information flow from the context in shallow attention layers, validating the effectiveness of Claim 1 and 2.
(a) Winogrande
(b) CSQA
Figure 4: Attention tracing results across different attention heads on Llama2-13B.
(a) Correct Case’s Attn
(b) Correct Case’s MLP
(c) Drifting Case’s Attn
(d) Drifting Case’s MLP
Figure 5: Intervention tracing results on Winogrande in correct and drifting answering cases.
3.3 Attention Tracing Experiment
Experimental Settings
We also design an experiment based on attention scores to validate Hypothesis 1 from another perspective. For a pair of rationales <$c,c^{*}$> targeting the same question context $q$ (c denotes the correct CoT and $c^{*}$ is the drifting one), we compute their attention divergence:
$Attn(c|c^{*})=\sum_{(i,j)\in C_{qc}}\frac{A^{(l,h)}_{i,j}}{|c|}-\sum_{(i,j)\in C_{qc^{*}}}\frac{A^{(l,h)}_{i,j}}{|c^{*}|}$ (4)
Here, we replace the $Attr(A_{i,j}^{(l)})$ in Equation 2 with the weights on the $h$-th attention matrix head $A_{i,j}^{(l,h)}$ and repeat the calculation in Equation $3$.
Result and Analysis
The results on two datasets are shown in Figure 4 and Figure 11. We can get the following observations: (1) The attention divergence is greater than 0 in most heads, which indicates a lack of information from the question context in Rationale Drift cases (consistent with Claim 1). (2) The largest attention divergence appears around layer 15, which is consistent with the sites we find in Claim 2. This once again illustrates that attention heads of these layers are crucial for the LLM to obtain contextual information when generating CoT.
4 Tracing Information Flow in Answer
[…]
4.1 Tracing Method
Since the task we study is in the form of multiple-choice questions, we set our focus on the feedforward pass that predicts the label. Inspired by the previous works (Meng et al., 2022; Stolfo et al., 2023; Geva et al., 2023), we take the causal tracing method to quantify the contribution of different intermediate variables during this pass. Specifically, for hidden states $h_{i}^{(l)}$ in a clean run that predicts the answer, we have:
$h_{i}^{(l)}=h_{i}^{(l-1)}+a_{i}^{(l)}+m_{i}^{(l)}$ (5)
$a_{i}^{(l)}=attn^{(l)}(h_{1}^{(l-1)},...,h_{i}^{(l-1)})$
$m_{i}^{(l)}=mlp^{(l)}(a_{i}^{(l)}+h_{i}^{(l-1)})$
where $i,l$ is the $i$-th token in the $l$-th layer, $a_{i}^{(l)},m_{i}^{(l)}$ represents the activations of attention and MLP modules in Transformer (Vaswani et al., 2017). Supposing that a certain input part is represented as $z=[v_{i}^{(l)},...,v_{j}^{(l)}]$ after passing through a model component, we set $v_{k}^{(l)}=v_{k}^{(l)}+\epsilon$ for $k\in[i,j]$ to intervene this hidden vector, where $\epsilon$ is Gaussian noise. [^4] We select $\epsilon$ to be 3 times larger than the empirical standard deviation of hidden embeddings in each dataset. Thus, we can compute the direct effect (DE) of this component:
$DE(z)=\frac{P(o)-P^{*}_{z}(o)}{P(o)}$ (6)
where $P(o)$ is the probability of the model’s final prediction, $P^{*}_{z}(o)$ is the probability after the intervention. Therefore, through this metric, we can quantify the contributions of different components in changing the final prediction, thereby tracing the information flow in this stage.
4.2 Intervention Tracing Experiment
Experimental Settings
We sample correct and drifting answering cases from datasets, average over them and compute the average direct effect (ADE). Here we compute the impact of four components on the final prediction: context (question contexts), option (question options), CoT, and last (the last token before the label prediction).
Results and Analysis
We report the result of Llama2-13B on Winogrande in Figure 5 and others in Appendix D, from which we can get two conclusions: (Claim 3) For attention modules, drifting cases loss information from the question. When the answer is correct, we can observe a high effect on the context and option in the first layer (see Figure 4(a)). But for the drifting case, the LLM extracts limited information at these positions (see Figure 4(c)). This aligns with Hypothesis 2. (Claim 4) For MLP modules, the information is not lost. We observe the same high-effect sites in the last layer and shallow layers of the last token, they do not show regular differences (see Figure 4(b) and 4(d)).
(a) Winogrande
(b) CSQA
Figure 6: Attribution tracing results on Llama2-13B during the answer generation stage.
4.3 Attribution Tracing Experiment
Experimental Settings
[…]
Results and Analysis
The results of this experiment are reported in Figure 6 and 15. We can observe that, when the Answer Drift issue occurs, the information flow from the question significantly decreases. This verifies the information loss we mention in Hypothesis 2 and Claim 3.
Figure 7: An example of our serial-position swap method.
5 Mitigating Toxic CoT Problem
[…]
5.1 Mitigation Method
Residual Decoding
We design a new decoding methodology to address the Rationale Drift issue, in which we construct a virtual residual structure during the CoT generation, “connecting” the question context with each CoT token. Our decoding algorithm is demonstrated in Algorithm 1. In each iteration of generating a new token, we first select the top $n$ tokens with the highest probabilities and record their logits scores (line 3). Then we calculate the attention score between the context and current token like Equation 4, normalize it, and add it as an additional reward to promote more information flow (lines 6,7). Finally, we select the token with the highest score to update the input and repeat the process until the termination condition is met. We use the attention matrix in layer 15 to compute the attention score, since it is crucial for the exchange of contextual information according to Claim 2. More implementation details of this method are provided in Appendix E.
Algorithm 1 Residual Decoding Algorithm
0:  model $\mathcal{M}$, input $x$, question context $q$, candidate_num $n$, weight $\omega$.
1:  for iteration $i\in 0,1,...$ do
2:     $logits=\mathcal{M}(x)$
3:     $tokens$, $scores$ = top($logits,n$)
4:     for $j\in 1,...n$ do
5:        $t=tokens[i]$
6:        $attn\_score=$ Attn($q$, $t$) / Attn($x$, $t$)
7:        $scores[j]=scores[j]$ + $\omega*attn\_score$
8:     end for
9:     $idx$ = argmax($scores$)
10:     $t=tokens[idx]$
11:     $x=x+t$
12:     if stop($t$then
13:        break
14:     end if
15:  end for
16:  return $x$
Serial-Position Swap
In this method, we attempt to compensate for the information lack in the Answer Drift issue. According to previous research on the serial-position effect in context, models tend to utilize information better at the beginning and end of the input (Qin et al., 2023; Liu et al., 2023b). In our work, the beginning of the input are prompts, while the current question and the generated CoT are both located at the end. Therefore, when they are closer to the last token, their information is more easily utilized in the final prediction. As in Figure 7, we denote the lengths of the question and CoT as $L_{q}$ and $L_{c}$, and assume that the key information is located at positions $\mu L_{q}$ and $\lambda L_{c}$ (similar to the center of mass in physics). We can infer that, after the swapping operation in Figure 7, the distance from the question to the end is reduced ($\mu L_{q}+L_{c}\rightarrow\mu L_{q}$). Besides, if we consider the total distance from the question and CoT to the end, we can perform the following calculation:
$d_{1}=\mu L_{q}+L_{c}+\lambda L_{c}$ (7)
$d_{2}=\lambda L_{c}+L_{q}+\mu L_{q}$
$\Delta d=d_{2}-d_{1}=L_{q}-L_{c}$
where $d_{1}$ is the total distance in normal serial positions and $d_{2}$ is the distance after swapping the two components. In most scenarios, we have $L_{q}<L_{c}$, thus, we can infer that $\Delta d<0$. That means, if we replace the original order of “[Question] + [CoT]” with the order of “[CoT] + [Question]”, we can not only increase the intensify of information flow from the question to the final prediction, but also reduce the total information loss due to the reduction in total distance. Although this method is straightforward in implementation, it proves to be effective in both theory and experiments.
Method Winogrande CSQA HellaSwag SIQA PIQA Avg
[…]
Few-shot Answer 57.1 - 66.8 - 45.3 - 68.2 - 61.6 - 59.8 -
Chain-of-Thought 56.7 43.1 69.9 28.8 45.4 34.8 67.3 37.3 62.7 41.3 60.4 37.1
Self-Consistency 56.4 40.9 72.4 24.9 46.2 33.7 64.7 41.0 54.6 45.8 58.9 37.3
Self-Refine 48.7 50.3 55.5 47.0 42.0 37.2 65.1 38.8 50.8 48.3 52.4 44.3
Least-to-Most 58.9 38.8 69.0 27.5 31.3 39.4 66.3 36.6 65.7 32.2 58.2 34.9
Contrasive CoT 59.4 38.0 71.1 23.2 45.9 35.6 67.0 36.0 68.1 44.9 62.3 35.5
Ours RD Only 58.6 29.5 72.2 13.5 49.4 24.0 69.4 20.8 66.3 26.2 63.0 22.8
SPS Only 59.1 22.8 72.6 18.9 49.3 18.7 69.7 28.8 68.5 27.6 63.7 23.4
RIDERS 60.7 16.6 73.2 6.7 50.6 12.8 71.9 15.3 69.9 15.9 65.3 13.5
Table 1: Performance comparison across five commonsense reasoning datasets on Llama2-13B.
5.2 Mitigation Experimental Settings
Datasets
Following previous works, we use five representative commonsense reasoning benchmarks: WinoGrande (Sakaguchi et al., 2020), CSQA (Talmor et al., 2019), HellaSwag (Zellers et al., 2019), SIQA (Sap et al., 2019) and PIQA (Bisk et al., 2020). The specific information of each dataset is reported in Appendix F.
Metrics
In addition to the commonly used Accuracy (ACC) metric, we also introduce a new metric — Toxic Rate (TR), to quantify the severity of Toxic CoT problems:
$TR(f)=|C_{d}\cap W_{f}|/|W_{f}|$ (8)
where $C_{d}$ denotes questions that models give correct answers directly and $W_{f}$ denotes questions that models give wrong answers after applying method $f$. Thus, we can infer that the lower the TR, the fewer Toxic CoT problems the method introduces.
Baselines
As our research focuses on enhancing CoT methods in commonsense reasoning, we select some of the latest CoT-like methods applicable to this task for comparison: Few-shot Answer, Chain-of-Thought (Wei et al., 2022), Self-Consistency (Wang et al., 2023c), Self-Refine (Madaan et al., 2023), Least-to-Most (Zhou et al., 2023) and Contrasive CoT (Chia et al., 2023). For all methods, we employ a 5-shot prompt and use 4 NVIDIA GeForce RTX 3090 GPUs for inference. More implementation details can be found in the Appendix F.
5.3 Mitigation Results
The main result of our experiments on Llama2-13B is shown in Table 1 (see Table 8 for the result on Baichuan2-13B). We can get the following conclusions: (1) Our method effectively mitigates Toxic CoT problems. Compared to CoT prompting, our method reduces the Toxic Rate by an average of 23.6% across five datasets. Besides, compared to other advanced CoT-like methods, our method causes the fewest Toxic CoT problems (decreased by an average of 22.0% over SOTA methods). (2) Our method can also improve the model’s overall performance on commonsense reasoning. Our work improves the accuracy on all benchmarks (improved by 5.5% compared to CoT and 3.0% compared to SOTA methods on average). This proves that the Toxic CoT problem poses a bottleneck in LLM’s commonsense reasoning, highlighting the value of our work.
Method Winogrande CSQA
Type1 Type2 Type1 Type2
CoT 0.0 0.0 0.0 0.0
RD Only 41.9 47.8 56.7 46.2
SPS Only 58.1 78.3 46.7 92.3
RIDERS 74.2 82.6 73.3 84.6
Table 2: Accuracy on the two types of drifting issues.
5.4 Discussion and Analysis
Performance on Two Drifting Issues
To demonstrate the effectiveness of our method in addressing the Rationale Drift issue (Type1) and Answer Drift issue (Type2), we conduct experiments on these samples and report the results in Table 2. Both of our methods can mitigate the corresponding issues (RD solves 49.3% Rationale Drift issue on average, while SPS resolves 85.3% Answer Drift issue on average). This verifies the validity of our hypothesis 1, 2, as all of these methods are built upon them. Besides, combining the two methods leads to even greater improvements, demonstrating the necessity of optimizing from both of these perspectives.
(a) RD method
(b) SPS method
Figure 8: Information flow comparison on Winogrande after applying our two methods.
Performance in the Model
[…]
Method Wino CSQA Hella SIQA PIQA Avg
CoT 3.2k 2.6k 4.2k 2.9k 2.7k 3.1k
SC 4.5k 3.2k 5.5k 3.8k 4.3k 4.3k
SR 9.2k 6.9k 10.2k 7.6k 6.5k 8.1k
L2M 8.7k 7.9k 11.4k 8.3k 6.7k 8.6k
CON 4.6k 3.5k 5.4k 3.8k 3.5k 4.2k
Ours 3.6k 2.9k 4.9k 3.4k 3.3k 3.6k
Table 3: Token consumption per example comparison.
Cost Analysis
For the applicability, we measure the computation and time cost of our approach. Here we compare the token cost between our method and the baseline. According to Table 3, our method requires fewer tokens compared to other SOTA methods (only 1.2×1.2$\times$1.2 × cost of the basic CoT method). We also compare the time cost of our decoding method in Appendix G and find that the speed of RD is comparable to existing decoding strategies. Therefore, we illustrate the cost-efficiency of our approach across different tasks.
6 Related Work
6.1 CoT Problems Analysis and Mitigation
Recently, many works have focused on analyzing and mitigating problems in CoT reasoning. For analytical work, most studies focus on black-box LLMs. Through intervening or paraphrasing prompts and comparing outputs, researchers can interpret the reasons leading to errors in the model’s reasoning (Lanham et al., 2023; Liu et al., 2023c; Wang et al., 2023a). For optimization works, they design additional supervision signals or training processes for the model (Ramnath et al., 2023; Liu et al., 2023a) or leverage external resources for the model (Shinn et al., 2023; He et al., 2023; Lyu et al., 2023). However, these works lack the probing of inner mechanisms behind these problems, leading to insufficient analysis or less universally applicable optimization methods.
6.2 Mechanistic Interpretability
The work on mechanistic interpretability aims to understand the internal mechanisms of models when performing various tasks. Early work focused on how the model stores factual knowledge internally (Meng et al., 2022; Dai et al., 2022). In recent times, some research efforts have shifted towards examining how models retrieve and utilize knowledge. This includes internal knowledge retrieval (Geva et al., 2023), knowledge retrieval from prompts (Wang et al., 2023b), and the utilization of knowledge for reasoning purposes, such as math reasoning (Stolfo et al., 2023) and multi-step reasoning (Hou et al., 2023). However, there is limited existing work that explains commonsense reasoning and CoT reasoning, which are significant contributions of this work.
7 Conclusion
In this paper, we find a problem named Toxic CoT, which results in the model’s reasoning deviating from the original correct answer when utilizing CoT-like prompting. Through tracing the internal information flow of the LLM with attribution tracing and causal tracing methods, we prove that this problem is mainly caused by the model’s lack of information from the question in shallow attention layers when generating rationales or answers. Based on this result, we propose the $\mathbb{RIDERS}$ method to mitigate the Toxic CoT problem from both decoding and serial position perspectives. Through extensive experiments on multiple commonsense reasoning datasets, we verify the effectiveness of our approach in mitigating Toxic CoT problems and enhancing the model’s overall commonsense reasoning capabilities.
Limitations
Although our work conducts an in-depth interpretation and mitigation of the Toxic CoT problem, it has several limitations. Firstly, like former commonsense reasoning works (Liu et al., 2022, 2023a; Xie et al., 2023), our research focuses on the form of multi-choice questions. This stems from the absence of effective evaluation methods for open-ended commonsense reasoning, leading to the predominance of benchmarks in this format. This calls for advancements in benchmark-related research. Secondly, we refrain from analyzing Toxic CoT problems in additional reasoning tasks such as math and logic, primarily due to the poor performance of current moderately-sized white-box models on these tasks. For instance, Llama2-13B achieves a mere 7.2% accuracy on GSM8K (Cobbe et al., 2021) without utilizing the CoT technique. This calls for developments in model-related research. We leave these limitations as our future work to explore.
Appendix A Early Statistical Experiments
In this section, we conduct early experiments on existing representative commonsense reasoning datasets to analyze the prevalence of Toxic CoT problems through statistical methods.
Datasets
We utilize five representative common-sense reasoning datasets to analyze the distribution of Toxic CoT problems. The basic information of the dataset is outlined in Table 4. It is noteworthy that, owing to the extensive size of Hellaswag’s dev set (over 10,000), we extract 2,000 instances for the experiment.
Metric
We design a new metric called Toxic Rate, which measures the proportion of Toxic CoT problems among all errors. Its calculation method is shown in Equation 8.
Results
The result of our early statistical experiments is reported in Table 5. Here we use Llama2-13B-Chat-hf to present the white-box LLM and use GPT-3.5-turbo-1106 to present the black-box-model. The average Toxic Rates are as high as 37.0% and 32.8% across the five datasets, indicating that this issue cannot be ignored and warrants further investigation.
Winogrande CSQA HellaSwag SIQA PIQA
Split dev dev dev dev dev
#Sample Num 1267 1221 2000 1954 1838
#Option Num 2 5 4 3 2
Table 4: Dataset information in this work.
Winogrande CSQA HellaSwag SIQA PIQA Avg
Llama2-13B 43.1 28.8 34.8 37.3 41.2 37.0
GPT-3.5 34.8 37.8 27.5 37.5 26.4 32.8
Table 5: Toxic rate on different datasets and models.
Appendix B Toxic Reason Statistical Experiments
In this section, we manually categorize the error types of Toxic CoT problems through statistical classification. Specifically, we sample 1,000 examples from CSQA and 1,000 examples from Winogrande, classifying the Toxic CoT problems (In all of the probing experiments in the main text, we use these samples as our probing data). The results are presented in Table 6. In the inconsistent error, the model exhibits logical inconsistency with the preceding context when generating CoT or the final answers. In the factual error, the CoT contains incorrect factual knowledge, which leads to erroneous answers. The presence of question errors reflects the subpar quality of the dataset. In such cases, questions may exhibit multiple viable answers or all options are incorrect. As for the other error, the questions trigger certain refusal-to-answer mechanisms in the model (e.g., inquiries about how to commit murder), leading to the identification of incorrect answers.
[…]
Reason Winogrande CSQA
Inconsistent Error 71(78.9%) 48(53.9%)
Factual Error 16(17.8%) 17(19.1%)
Question Error 2(2.2%) 21(23.6%)
Other Error 1(1.1%) 3(3.4%)
Sum 90 89
Table 6: The classification of CoT reasoning errors
Appendix C More Details for Reasoning Tracing
Method Implementation
[…]
Attribution Tracing Experiment
Strategy Wino CSQA Hella SIQA PIQA Avg
Greedy 5.2 2.8 4.9 3.6 5.8 4.5
Beam 8.1 4.9 10.1 5.8 9.4 7.7
Ours 13.4 6.6 16.2 7.4 15.9 11.9
Table 7: Decoding time per example comparison.
Figure 17 illustrates the prompt we use for generating correct CoTs from drifting cases. After the generation, we will manually filter out the wrong CoT and conduct the comparative experiment. We use a 5-shot to generate CoT and concatenate it to the question for our probing experiments. Figure 9 shows the remaining results of this experiment, from which we can get the same conclusions as Section 3.2. Additionally, we also conduct the supplementary experiments on CSQA and report the results in Figure 10.
Attention Tracing Experiment
Figure 11 reports more results in this experiment.
Appendix D More Details for Answering Tracing
Intervention Tracing Experiment
Figure 12, 13, 14 and 15 report the remaining intervention tracing experiment results, which are consistent with our conclusion in the main text.
Appendix E Mitigation Method Implementation
Residual Decoding
Here, we provide a detailed explanation of Algorithm 1. At the beginning, we set the input to the entire question (contexts + options). In line 2, we get the logits from the output of the LlamaForCausalLM. In line 6, we calculate the attention score by summing the values on the attention matrix corresponding to the tokens. We use the output character “</s>” as the termination condition for Llama2-13B generation in line 12.
Serial-Position Swap
In this method, we swap the positions of the question and the generated CoT, outputting the option with the highest logits score. This method can be implemented under both few-shot and zero-shot settings, demonstrating its cost-efficiency.
Appendix F More Details for the Mitigation Experiment
Dataset
In this experiment, the specific information of all datasets can be found in Table 4.
Baselines
For the Self-Consistency method, we sample 5 CoTs and use a majority voting method to select the final predicted answer. For the Self-Refine method, we first conduct one round of CoT reasoning and then follow it with one round of feedback to generate the final answer. For all the baselines, we will release the prompts in our source code if the work is accepted. We implement all of the methods on Llama2-13B-Chat-hf and Baichuan2-13B-Chat.
Our methods
In our RD method, we set two hyperparameters — candidate_num $n$ and weight $\omega$, and here are their specific values in the experiments: for Winogrande, we set $n$ to 4 and $\omega$ to 80, for CSQA, we set $n$ to 10 and $\omega$ to 135, for HellaSwag, we set $n$ to 3 and $\omega$ to 80, for SIQA, we set $n$ to 10 and $\omega$ to 160, for PIQA, we set $n$ to 4 and $\omega$ to 120. Additionally, in Figure 18,19,20,21,22, we list our method’s few-shot prompts on five datasets. Note that both CoT prompting and our two methods utilize the same prompt.
Results
Table 8 reports more results in this experiment.
Appendix G More Details for Cost Analysis
Table 7 illustrates the time cost comparison of our residual decoding methods with other strategies. Here we set num_beams in the beam search strategy as 5, and candidate_num in the RD strategy as 3. We compute the average seconds cost per example over 50 samples for each dataset. On average, our decoding strategy takes 2.6 times longer than greedy search and 1.4 times longer than beam search. This reflects that our decoding method has significantly stronger performance while having a comparable overall time to these main decoding strategies.
Method Winogrande CSQA HellaSwag SIQA PIQA Avg
[…]
Few-shot Answer 59.7 - 69.9 - 47.1 - 68.8 - 60.7 - 61.2 -
Chain-of-Thought 58.7 34.8 66.8 37.8 41.1 27.5 66.5 37.5 66.1 26.4 59.8 32.8
Self-Consistency 56.1 38.8 64.3 41.5 39.4 30.6 67.8 33.5 70.0 21.0 59.5 33.1
Self-Refine 59.6 35.7 66.7 38.8 39.2 29.9 62.2 43.1 62.6 31.0 58.1 35.7
Least-to-Most 59.7 35.6 62.8 45.6 37.8 32.1 66.5 36.8 62.8 20.6 57.9 34.1
Contrasive CoT 55.2 40.6 67.0 37.4 38.7 29.7 65.0 37.3 65.9 26.2 58.4 34.2
Ours RD Only 59.2 21.1 69.5 25.3 44.4 19.3 68.9 25.0 67.0 12.9 61.8 20.7
SPS Only 59.9 22.8 69.7 23.5 46.3 15.2 69.2 27.8 67.7 14.2 62.6 20.7
RIDERS 60.1 18.6 71.3 14.6 46.3 13.0 69.0 20.5 67.2 8.3 62.8 15.0
Table 8: Performance comparison across five commonsense reasoning datasets on Baichuan2-13B.
(a) Llama2-13B
(b) Baichuan2-13B
Figure 9: Attribution tracing results on CSQA.
(a) Llama2-13B
(b) Baichuan2-13B
Figure 10: Information flow divergence comparison on CSQA.
(a) Winogrande
(b) CSQA
Figure 11: Attention tracing results across different attention heads on Baichuan2-13B.
(a) Correct Case’s Attn
(b) Correct Case’s MLP
(c) Drifting Case’s Attn
(d) Drifting Case’s MLP
Figure 12: Intervention tracing results on CSQA in correct and drifting answering cases (Llama2-13B).
(a) Correct Case’s Attn
(b) Correct Case’s MLP
(c) Drifting Case’s Attn
(d) Drifting Case’s MLP
Figure 13: Intervention tracing results on Winogrande in correct and drifting answering cases (Baichuan2-13B).
(a) Correct Case’s Attn
(b) Correct Case’s MLP
(c) Drifting Case’s Attn
(d) Drifting Case’s MLP
Figure 14: Intervention tracing results on CSQA in correct and drifting answering cases (Baichuan2-13B).
(a) Winogrande
(b) CSQA
Figure 15: Attribution tracing results on Baichuan2-13B during the answer generation stage.
(a) Winogrande
(b) CSQA
Figure 16: Information flow comparison on CSQA after applying our two methods.
Figure 17: Prompts for correct CoT generation.
Figure 18: 5-shot prompts for Winogrande.
Figure 19: 5-shot prompts for CSQA.
Figure 20: 5-shot prompts for HellaSwag.
Figure 21: 5-shot prompts for SIQA.
Figure 22: 5-shot prompts for PIQA.

### v2

@@
 Through extensive experiments on multiple commonsense reasoning benchmarks, we validate that this method not only significantly eliminates Toxic CoT problems (decreased by 23.6%), but also effectively improves the model’s overall commonsense reasoning performance (increased by 5.5%).
-Focus on Your Question! Interpreting and Mitigating Toxic CoT Problems in Commonsense Reasoning
-Jiachun Li 1,2, Pengfei Cao 1,2, Chenhao Wang 1,2, Zhuoran Jin 1,2, Yubo Chen 1,2 Daojian Zeng 3, Kang Liu 1,2, Jun Zhao 1,2 1The Laboratory of Cognition and Decision Intelligence for Complex Systems, Institute of Automation, Chinese Academy of Sciences 2School of Artificial Intelligence, University of Chinese Academy of Sciences 3Hunan Normal University {jiachun.li, pengfei.cao, chenhao.wang, zhuoran.jin, yubo.chen, kliu, jzhao} @nlpr.ia.ac.cn zengdj916@163.com
+Focus on Your Question! Interpreting and Mitigating Toxic CoT
+Problems in Commonsense Reasoning
+Jiachun Li1,2, Pengfei Cao1,2, Chenhao Wang1,2, Zhuoran Jin1,2, Yubo Chen1,2, [^1] Corresponding authors.
+Daojian Zeng3, Kang Liu1,2, Jun Zhao1,2, [^1] Corresponding authors.
+1School of Artificial Intelligence, University of Chinese Academy of Sciences 2The Laboratory of Cognition and Decision Intelligence for Complex Systems, Institute of Automation, Chinese Academy of Sciences 3Hunan Normal University {jiachun.li, pengfei.cao, chenhao.wang, zhuoran.jin, yubo.chen, kliu, jzhao} @nlpr.ia.ac.cn zengdj916@163.com
 1 Introduction
-With the increase in scale, large language models (LLMs) have demonstrated outstanding performance in different reasoning tasks, among them, commonsense reasoning has received significant attention due to its importance for general intelligence.
+With the increase in scale, large language models (LLMs) have demonstrated outstanding performance in different tasks (Li et al., 2023; Wen et al., 2023; Sun et al., 2024; Jin et al., 2024a), among them, commonsense reasoning has received significant attention due to its importance for general intelligence (Wang et al., 2022, 2024; Liu et al., 2024).
 In this task, researchers have proposed a series of chain-of-thought (CoT) like techniques to elicit models’ potential abilities (e.g.
@@
 While these works have made great progress, we notice an overlooked problem in them, which we define as Toxic CoT — Sometimes LLMs can directly provide correct answers to questions, but after applying CoT-like methods, it brings extra reasoning paths to models, causing their answers to be wrong.
+[^1] Notably, our definition here differs from the toxicity in generated content (Shaikh et al., 2023), emphasizing the harm that CoT can bring to the model.
 Figure 1 illustrates two main error types of this problem — Rationale Drift and Answer Drift.
@@
 The results not only verify the rationality of our interpretation, but also demonstrate the effectiveness of our method in addressing the Toxic CoT problem (the proportion of the problem decreased by 23.6%) and enhancing commonsense reasoning performance (overall accuracy increased by 5.5%).
-We will release the source code for this work if it is accepted.
+Our code is available at: https://github.com/BugMakerzzz/toxic_cot.
 2 Problem Statement
@@
 3 Tracing Information Flow in Rationale
-[…]
+In this section, we aim to verify the Hypothesis 1 by tracing the information flow in the rationale generation stage.
+To this end, we start by describing our attribution tracing method (§$\S$3.1).
+Through this method, we conduct comparative experiments between the correct reasoning and the drifting one, figuring out the mechanism behind the issue (§$\S$3.2).
+At last, we use the attention score to validate our findings from another perspective (§$\S$3.3).
 3.1 Tracing Method
@@
 4 Tracing Information Flow in Answer
-[…]
+In this section, our goal is to verify the information loss based on Hypothesis 2.
+To achieve this, we first introduce the main tracing method in this section, which is called the causal tracing method (§$\S$4.1).
+Next, by employing it, we trace the information flow in the answer generation stage and identify the mechanism behind the Answer Drift issue through comparative experiments (§$\S$4.2).
+At last, we apply the attribution tracing method to verify our hypothesis from another perspective (§$\S$4.3).
 4.1 Tracing Method
@@
 (Claim 4) For MLP modules, the information is not lost.
-We observe the same high-effect sites in the last layer and shallow layers of the last token, they do not show regular differences (see Figure 4(b) and 4(d)).
+We observe the same high-effect sites in the last layer and shallow layers of the last token, they do not show regular differences (see Figure 5(b) and 5(d)).
 (a) Winogrande
@@
 Experimental Settings
-[…]
+For further validation of our hypothesis, we also use the attribution score in §$\S$3 to trace the information flow in this stage.
+Referring to Equation 2, we compute the score between the question context and the last token (since it’s used for generating the answer).
+We set the $F(\cdot)$ in Equation 1 as the loss for predicting the final answer, comparing the scores for correct and drifting cases after averaging across samples.
 Results and Analysis
@@
 5 Mitigating Toxic CoT Problem
-[…]
+In this section, we propose a novel method called $\mathbb{RIDERS}$ (Residual decodIng and sERial-position Swap) to address the Toxic CoT problem.
+We first introduce the two components in it, which are designed based on Hypothesis 1 and 2, respectively (§$\S$5.1).
+Then, we conduct experiments on commonsense reasoning benchmarks, demonstrating the effectiveness of our method (§$\S$5.2 and §$\S$5.3).
+At last, we conduct extra experiments to further emphasize the contribution of our approach (§$\S$5.4).
 5.1 Mitigation Method
@@
 5.3 Mitigation Results
-The main result of our experiments on Llama2-13B is shown in Table 1 (see Table 8 for the result on Baichuan2-13B).
-We can get the following conclusions: (1) Our method effectively mitigates Toxic CoT problems.
+The main result of our experiments on Llama2-13B is shown in Table 1.
+(The results on more models are presented in Appendix F.) We can get the following conclusions: (1) Our method effectively mitigates Toxic CoT problems.
 Compared to CoT prompting, our method reduces the Toxic Rate by an average of 23.6% across five datasets.
@@
 Performance in the Model
+In §$\S$3 and §$\S$4, we probe information loss in two issues by tracing the information flow in models.
+Here, we repeat the attribution tracing experiments, comparing the differences before and after applying our method to further validate the effectiveness of our work.
+As we can see from Figure 8 and 16, our two methods (orange lines) increase the information flow from questions in two stages compared to CoT prompting (blue lines).
+This indicates our method indeed compensates for the information loss in the LLM.
+Method ProofWriter FOLIO
 […]
+CoT 56.8 19.9 45.1 11.6
+RIDERS 78.2 6.4 48.0 7.5
+Table 3: Performance comparison across two logical reasoning tasks on Llama2-13B.
+Performance on Other Tasks
+Our work mainly focuses on commonsense reasoning tasks.
+To further evaluate the effectiveness of our approach on other types and forms of tasks, we conduct the main experiments on two logical reasoning tasks: ProofWriter (Tafjord et al., 2021) and FOLIO (Han et al., 2022)).
+As presented in Table 3, we can find that the Toxic CoT problem also exists in other tasks, and our method can still mitigate it.
 Method Wino CSQA Hella SIQA PIQA Avg
@@
 In recent times, some research efforts have shifted towards examining how models retrieve and utilize knowledge.
-This includes internal knowledge retrieval (Geva et al., 2023), knowledge retrieval from prompts (Wang et al., 2023b), and the utilization of knowledge for reasoning purposes, such as math reasoning (Stolfo et al., 2023) and multi-step reasoning (Hou et al., 2023).
+This includes internal knowledge retrieval (Geva et al., 2023), knowledge retrieval from prompts (Wang et al., 2023b; Jin et al., 2024b), and the utilization of knowledge for reasoning purposes, such as math reasoning (Stolfo et al., 2023) and multi-step reasoning (Hou et al., 2023).
 However, there is limited existing work that explains commonsense reasoning and CoT reasoning, which are significant contributions of this work.
@@
 This calls for advancements in benchmark-related research.
-Secondly, we refrain from analyzing Toxic CoT problems in additional reasoning tasks such as math and logic, primarily due to the poor performance of current moderately-sized white-box models on these tasks.
+Secondly, we refrain from analyzing Toxic CoT problems in more reasoning tasks such as math, primarily due to the poor performance of current moderately-sized white-box models on these tasks.
 For instance, Llama2-13B achieves a mere 7.2% accuracy on GSM8K (Cobbe et al., 2021) without utilizing the CoT technique.
@@
 We leave these limitations as our future work to explore.
+Acknowledgement
+This work is supported by the Strategic Priority Research Program of Chinese Academy of Sciences (No. XDA27020203), the National Natural Science Foundation of China (No. 62176257,62276095).
 Appendix A Early Statistical Experiments
@@
 As for the other error, the questions trigger certain refusal-to-answer mechanisms in the model (e.g., inquiries about how to commit murder), leading to the identification of incorrect answers.
-[…]
+As the inconsistency error constitutes the predominant portion of all reasons, our work focuses on addressing this issue.
+We further categorize this error into Rationale Drift and Answer Drift based on the error occurrence (see §$\S$2.2 for their definitions).
 Reason Winogrande CSQA
@@
 Method Implementation
-[…]
+We introduce the attribution score method in §$\S$ 3.1.
+In Equation 1, we set $m$ = 20 following the previous works.
+For $F(\cdot)$, we set it as the language modeling loss (for next-token prediction) during the CoT generation.
+Here, we obtain this value directly from the output of the LlamaForCausalLM module using the Transformers library.
+In Equation 2, we partition the step numbers $|N|$ in CoT based on the occurrence of periods in the text.
+For models, we use Llama2-13B-Chat and Baichuan2-13B-Chat.
 Attribution Tracing Experiment
-Strategy Wino CSQA Hella SIQA PIQA Avg
-Greedy 5.2 2.8 4.9 3.6 5.8 4.5
-Beam 8.1 4.9 10.1 5.8 9.4 7.7
-Ours 13.4 6.6 16.2 7.4 15.9 11.9
-Table 7: Decoding time per example comparison.
 Figure 17 illustrates the prompt we use for generating correct CoTs from drifting cases.
@@
 For the Self-Refine method, we first conduct one round of CoT reasoning and then follow it with one round of feedback to generate the final answer.
-For all the baselines, we will release the prompts in our source code if the work is accepted.
+For all the baselines, we release the prompts in our source code.
 We implement all of the methods on Llama2-13B-Chat-hf and Baichuan2-13B-Chat.
@@
 Note that both CoT prompting and our two methods utilize the same prompt.
-Results
-Table 8 reports more results in this experiment.
-Appendix G More Details for Cost Analysis
-Table 7 illustrates the time cost comparison of our residual decoding methods with other strategies.
-Here we set num_beams in the beam search strategy as 5, and candidate_num in the RD strategy as 3.
-We compute the average seconds cost per example over 50 samples for each dataset.
-On average, our decoding strategy takes 2.6 times longer than greedy search and 1.4 times longer than beam search.
-This reflects that our decoding method has significantly stronger performance while having a comparable overall time to these main decoding strategies.
 Method Winogrande CSQA HellaSwag SIQA PIQA Avg
@@
 Table 8: Performance comparison across five commonsense reasoning datasets on Baichuan2-13B.
+Model Method Winogrande SIQA
+[…]
+Mistral-7B CoT 59.2 27.9 70.2 28.2
+RIDERS 62.0 15.3 70.6 24.5
+GPT-3.5 CoT 64.6 48.8 70.6 30.3
+RIDERS 70.1 41.7 72.1 29.0
+Table 9: Performance comparison on Mistral-7B and GPT-3.5.
+Results
+In our paper, we choose Llama2-13B for experiments because it is an open-source LLM of suitable size and has great influence.
+Here, we also repeat all experiments on Baichuan2-13B to verify the generality of our work (see Table 8).
+Additionally, in Table 9, we show our experimental results on two other representative models: Mistral-7B-Instruct-v0.2 and gpt-3.5-turbo-1106 (since it is a closed source model, here we only apply the SPS method).
+Strategy Wino CSQA Hella SIQA PIQA Avg
+Greedy 5.2 2.8 4.9 3.6 5.8 4.5
+Beam 8.1 4.9 10.1 5.8 9.4 7.7
+Ours 13.4 6.6 16.2 7.4 15.9 11.9
+Table 10: Decoding time per example comparison.
+Appendix G More Details for Cost Analysis
+Table 10 illustrates the time cost comparison of our residual decoding methods with other strategies.
+Here we set num_beams in the beam search strategy as $5$, and candidate_num in the RD strategy as $3$.
+We compute the average seconds cost per example over 50 samples for each dataset.
+On average, our decoding strategy takes 2.6 times longer than greedy search and 1.4 times longer than beam search.
+This reflects that our decoding method has significantly stronger performance while having a comparable overall time to these main decoding strategies.
 (a) Llama2-13B